What is the specific heat of a metal sample if a 35 g sample absorbs 2.3 x 10-3 MJ of heat and its temperature changes from 25°C to 175°C?



1
Expert's answer
2021-01-19T04:39:47-0500

The specific heat capacity of a substance is the amount of heat required to raise the temperature of 1g of the substance by 1oC. First, we need to convert the units of the quantities into workable units;

From Q = mC"\Delta"T

m = 35g

m=?

Q=(2.3 x 10-3MJ) x 1 000 000 J = 2300J

"\Delta"T = (175 - 25)oC = 150oC


Solve for C in the formula;


C "=\dfrac{C}{m\Delta T } = \dfrac{2300J}{35g x 150^oC} = 0.4381" Jg-1oC-1